Prof. Dr. M. Kreuzer Abgabe: Briefkasten 63, 3. Februar 2004, 12:00 Uhr

Algebra I

Übungsblatt 14

Through the laws of comparative linguistics one can trace a word like eleven to one left over, and similarly twelve to two over. There is some indication that our fundamental word ten may be derived from an Indo-European root meaning two hands. The word hundred comes from an original term ten times (ten). It is further interesting to note that the names for thousand are unrelated in the various main branches of the Indo-European languages; hence it is probably a rather late construction.

[…]

Also, in many languages one finds other methods for the construction of numbers. As an example of irregular construction let us mention that in Welsh the number words from 15 to 19 indicate 15, 15+1, 15+2, 2×9, 15+4.

Oystein Ore: Number Theory and Its History

Aufgabe 1: Sei p eine Primzahl der Form 2^{m} + 1 (Fermat-Primzahl). Zeigen Sie, m = 2^{k} für ein geeignetes k.

Aufgabe 2: Berechnen Sie die Kreisteilungspolynome Φ_{n} für n = 1, 2, …, 8.

Aufgabe 3: Sei ζ eine primitive zwölfte Einheitswurzel. Berechnen Sie die Galoisgruppe der Erweiterung ℚ(ζ)|ℚ. Bestimmen Sie alle Zwischenkörper der Erweiterung.

Aufgabe 4: Sei p eine Primzahl und ζ eine primitive p-te Einheitswurzel. Zeigen Sie, dass ζ, ζ^{2}, …, ζ^{p−1} eine ℚ-Basis der Erweiterung ℚ(ζ)|ℚ bilden.